OpenStudy (anonymous):

PLEASE HELP? Given that DEFG is a square, find x and y. A. x = 16, y = 9 B. x = 27, y = 4 C. x = 7, y = 9 D. not enough information

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

hint : a square has all right angles

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

grab the "y" equations and make them equal, because both legs are equal, then solve for "y"

OpenStudy (anonymous):

5y-4=2y+23

OpenStudy (jdoe0001):

get ganeshie8 's hint and make that equal to what a right-angle is :), then solve for "x"

OpenStudy (anonymous):

5y-4=90 2y+23=90

OpenStudy (anonymous):

like that?

OpenStudy (jdoe0001):

5y-4 = 2y+23

ganeshie8 (ganeshie8):

i thought so u would do that lol

OpenStudy (anonymous):

hahah.x) so i was wrong?

OpenStudy (jdoe0001):

left length = right length

OpenStudy (anonymous):

confused..

OpenStudy (jdoe0001):

"x equation" degrees = "right-angle" degreess

OpenStudy (anonymous):

so its like this? 5y-4 = 2y+23

OpenStudy (jdoe0001):

look at your picture, is the left-line equal to the right-line?

OpenStudy (anonymous):

ok?

OpenStudy (jdoe0001):

is a square, so all sides are equal, so

OpenStudy (jdoe0001):

the equation are just the length

OpenStudy (anonymous):

hmm..?

OpenStudy (jdoe0001):

so, if both lines are equal, so are their equations

OpenStudy (anonymous):

D: Your confusing me.

OpenStudy (jdoe0001):

I wish I had 2 chicken to use for comparison but anyway

OpenStudy (anonymous):

draw a picture??